# Mirror and Preconditioned Gradient Descent in Wasserstein Space

### v1

@@ D.1 Optimal transport maps for mirror descent
Let $\phi:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a strictly convex functional along all acceleration-free curves and denote for $\mu\in L^{2}(\mu)$, $\phi_{\mu}(\mathrm{T})=\phi(\mathrm{T}_{\#}\mu)$. Since $\phi$ is strictly convex along all acceleration-free curves, by Proposition 11, for all $\mathrm{T}\neq\mathrm{S}\in L^{2}(\mu)$, $\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S})>0$ and thus $\phi_{\mu}$ is strictly convex. Recall that
@@
1. Observe that problem (43) is equivalent with
@@
We continue this section with additional results relative to the invertibility of mirror maps, which are required in Proposition 1.
@@
Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and let $W:\mathbb{R}^{d}\to\mathbb{R}$ be even, $\epsilon$-strongly convex for $\epsilon>0$ and differentiable, then $\nabla_{\mathrm{W}_{2}}\mathcal{W}(\mu)$ is invertible.
@@
On one hand, $\nabla_{\mathrm{W}_{2}}\mathcal{W}(\mu)=\nabla W\star\mu$. Moreover, $W$ $\epsilon$-strongly convex is equivalent with
@@
Let $\mu\in\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$ such that its density is of the form $\rho\propto e^{-V}$ with $V:\mathbb{R}^{d}\to\mathbb{R}$ $\epsilon$-strongly convex for $\epsilon>0$, then $\nabla_{\mathrm{W}_{2}}\mathcal{H}(\mu)$ is invertible.
@@
The OT problem (43) recovers other OT costs for specific choices of $\phi$. For instance, for $\phi_{\mu}(\mathrm{T})=\frac{1}{2}\|\mathrm{T}\|_{L^{2}(\mu)}^{2}$, it coincides with the squared 2-Wasserstein distance. And more generally, for $\phi_{\mu}^{V}(\mathrm{T})=\int V\circ\mathrm{T}\ \mathrm{d}\mu$, since by Lemma 29, for all $\mathrm{T},\mathrm{S}\in L^{2}(\mu)$,
@@ D.2 Continuous formulation
Let $\phi:L^{2}(\mu)\to\mathbb{R}$ be pushforward compatible and superlinear. Introducing the (mirror) map $\varphi(\mu)=\nabla_{\mathrm{W}_{2}}\phi(\mu)$, we can write informally the discrete scheme (4) and its limit when $\tau\to 0$ as
@@
However, $\frac{\mathrm{d}}{\mathrm{d}t}\varphi(\mu_{t})=\frac{\mathrm{d}}{\mathrm{d}t}\nabla_{\mathrm{W}_{2}}\phi(\mu_{t})=\mathrm{H}\phi_{\mu_{t}}(v_{t})$ where $\mathrm{H}\phi_{\mu_{t}}:L^{2}(\mu_{t})\to L^{2}(\mu_{t})$ is the Hessian operator defined such that $\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\phi(\mu_{t})=\langle\mathrm{H}\phi_{\mu_{t}}(v_{t}),v_{t}\rangle_{L^{2}(\mu_{t})}$ and $v_{t}\in L^{2}(\mu_{t})$ is a velocity field satisfying $\partial_{t}\mu_{t}+\mathrm{div}(\mu_{t}v_{t})=0$. Thus, the continuity equation followed by the Mirror Flow is given by
@@ D.3 Derivation in specific settings
In this Section, we analyze several new mirror schemes obtained through different Bregman potential maps. We start by discussing the scheme with an interaction energy as Bregman potential. Then, we study mirror descent with negative entropy or KL divergence as Bregman potential. For the last two, we derive closed-forms for the case where every distribution is Gaussian, which is equivalent with working on the Bures-Wasserstein space, and to use the gradient on the Bures-Wasserstein space [40]. In particular, this space is a submanifold of $\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$ and the tangent space is the space of affine maps with symmetric linear term, i.e. of the form $T(x)=b+S(x-m)$ with $S\in S_{d}(\mathbb{R})$.
@@
Let us take as Bregman potential $\phi_{\mu}(\mathrm{T})=\iint W\big{(}(T(x)-T(x^{\prime})\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$. The general scheme is given by
@@
For the particular case $W(x)=\frac{1}{2}\|x\|_{2}^{2}$, the scheme can be made more explicit as $\nabla W\star\mu(x)=\int\nabla W(x-y)\ \mathrm{d}\mu(y)=\int(x-y)\ \mathrm{d}\mu(y)=x-m(\mu)$ with $m(\mu)=\int y\mathrm{d}\mu(y)$ the expectation, and thus it translates as
$\forall k\geq 0,\ x_{k+1}-m(\mu_{k+1})=x_{k}-m(\mu_{k})-\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}).$ (55)
@@
since $\nabla^{2}W=I_{d}$. On the other hand, the scheme can be written as, for all $k\geq 0$,
$\nabla_{\mathrm{W}_{2}}\phi(\mu_{k+1})(x_{k+1})=\nabla_{\mathrm{W}_{2}}\phi(\mu_{k})(x_{k})-\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})(x_{k})$ (57)
$\iff x_{k+1}-m(\mu_{k+1})=x_{k}-m(\mu_{k})-\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})(x_{k})$
$\iff\begin{cases}y_{k}=x_{k}-m(\mu_{k})\\ y_{k+1}=y_{k}-\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})(x_{k}).\end{cases}$
@@
$\begin{cases}y_{t}=x_{t}-m(\mu_{t})\\ \frac{\mathrm{d}y_{t}}{\mathrm{d}t}=-\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{t})(x_{t}).\end{cases}$ (58)
But $\frac{\mathrm{d}y_{t}}{\mathrm{d}t}=\frac{\mathrm{d}x_{t}}{\mathrm{d}t}-\frac{\mathrm{d}m(\mu_{t})}{\mathrm{d}t}$. Now, by setting $v_{t}(x)=\frac{\mathrm{d}x_{t}}{\mathrm{d}t}$ and noting that, by integration by part,
$\frac{\mathrm{d}}{\mathrm{d}t}m(\mu_{t})=\int x\ \partial_{t}\mu_{t}=-\int x\cdot\mathrm{div}(\mu_{t}v_{t})=\int v_{t}(y)\ \mathrm{d}\mu_{t}(y),$ (59)
we obtain indeed $\frac{\mathrm{d}y_{t}}{\mathrm{d}t}=\mathrm{H}\phi_{\mu_{t}}[v_{t}](x)$.
@@
Let us consider $\phi(\mu)=\int\log\big(\rho(x)\big{missing})\ \mathrm{d}\mu(x)$ where $\mathrm{d}\mu(x)=\rho(x)\mathrm{d}x$. For such Bregman potential, the mirror scheme can be obtained for all $k\geq 0$ as
@@
In general, this scheme is not tractable. Nonetheless, supposing that $\mu_{k}=\mathcal{N}(m_{k},\Sigma_{k})$ for all $k$, the scheme translates as
$-\Sigma_{k+1}^{-1}(\mathrm{T}_{k+1}(x)-m_{k+1})=-\Sigma_{k}^{-1}(x-m_{k})-\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}).$ (61)
For $\mathcal{F}(\mu)=\mathcal{H}(\mu)+\mathcal{V}(\mu)$ with $V(x)=\frac{1}{2}x^{T}\Sigma^{-1}x$, the scheme is
@@
Assuming $m_{k}=0$ for all $k$ and taking the covariance, then we obtain the following update rule for the covariance matrices:
@@
For $\mathcal{F}(\mu)=\mathcal{H}(\mu)$, we obtain
$-\Sigma_{k+1}^{-1}(\mathrm{T}_{k+1}(x)-m_{k+1})=-(1-\tau)\Sigma_{k}^{-1}(x-m_{k}).$ (64)
Assuming $m_{k}=0$ for all $k$, for $\tau<1$, taking the covariance, we get
$\Sigma_{k+1}^{-1}\Sigma_{k+1}\Sigma_{k+1}^{-1}=(1-\tau)^{2}\Sigma_{k}^{-1}\Sigma_{k}\Sigma_{k}^{-1}\\ \iff\Sigma_{k+1}=\frac{1}{(1-\tau)^{2}}\Sigma_{k}=\frac{1}{(1-\tau)^{2k}}\Sigma_{0}\underset{\tau\to 0}{\sim}e^{2\tau k}\Sigma_{0}.$ (65)
The underlying flow is thus $t\mapsto\mathcal{N}(0,e^{2t}\Sigma_{0})$ and the negative entropy decreases along this curve as
$\mathcal{H}(\rho_{t})=-\frac{d}{2}\log(2\pi e)-dt-\sum_{i=1}^{d}\log(\lambda_{i}),$ (66)
where $(\lambda_{i})_{i}$ denote the eigenvalues of $\Sigma_{0}$. This is much faster than the heat flow for which the negative entropy decreases as [118, Appendix E.2]
@@
with the scheme given by [118, Example 6]
@@
When we want to optimize the KL divergence, i.e. a functional of the form $\mathcal{F}(\mu)=\mathcal{H}(\mu)+\int V\mathrm{d}\mu$, then a natural choice of Bregman potential is also a functional of the form $\phi(\mu)=\mathcal{H}(\mu)+\int\psi\mathrm{d}\mu$ with $V$ $\alpha$-convex and $\beta$-smooth relative to $\psi$. Indeed, usually, for mirror maps only composed from a potential, the non smooth term is $\mathcal{H}$.
Let us note $\mathcal{F}(\mu)=\mathcal{G}(\mu)+\mathcal{H}(\mu)$ where $\mathcal{G}(\mu)=\int V\mathrm{d}\mu$ and $\phi(\mu)=\Psi(\mu)+\mathcal{H}(\mu)$ with $\Psi(\mu)=\int\psi\mathrm{d}\mu$. In that case, we have, since $\mathrm{d}_{\tilde{\mathcal{H}}_{\mu}+\tilde{\mathcal{G}}_{\mu}}=\mathrm{d}_{\tilde{\mathcal{H}}_{\mu}}+\mathrm{d}_{\tilde{\mathcal{G}}_{\mu}}$, for all $\mathrm{T},\mathrm{S}\in L^{2}(\mu)$,
$\mathrm{d}_{\tilde{\mathcal{F}}_{\mu}}(\mathrm{T},\mathrm{S})$ $=\mathrm{d}_{\tilde{\mathcal{H}}_{\mu}}(\mathrm{T},\mathrm{S})+\mathrm{d}_{\tilde{\mathcal{G}}_{\mu}}(\mathrm{T},\mathrm{S})$ (69)
$\leq\mathrm{d}_{\tilde{\mathcal{H}}_{\mu}}(\mathrm{T},\mathrm{S})+\beta\mathrm{d}_{\tilde{\Psi}_{\mu}}(\mathrm{T},\mathrm{S})$
$\leq\mathrm{max}(1,\beta)\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S}).$
@@
We now focus on the case where all measures are Gaussian in order to be able to compute a closed-form, i.e. $V(x)=\frac{1}{2}(x-m)^{T}\Sigma^{-1}(x-m)$, $\psi(x)=\frac{1}{2}x^{T}\Lambda^{-1}x$ and for all $k$, $\mu_{k}=\mathcal{N}(m_{k},\Sigma_{k})$. In this case, remember that $\nabla\log\mu_{k}(x)=-\Sigma_{k}^{-1}(x-m_{k})$. Then, at each step, the scheme is
$\nabla V(x_{k+1})+\nabla\log\big(\mu_{k+1}(x_{k+1})\big{missing})=\nabla V(x_{k})+\nabla\log\big(\mu_{k}(x_{k})\big{missing})-\tau\big{(}\nabla U(x_{k})+\nabla\log\big(\mu_{k}(x_{k})\big{missing})$ (70)
@@
$\quad\quad\quad=\big{(}\Lambda^{-1}-(1-\tau)\Sigma_{k}^{-1}-\tau\Sigma^{-1}\big{)}x_{k}+(1-\tau)\Sigma_{k}^{-1}m_{k}+\tau\Sigma^{-1}m.$
@@
$(\Lambda^{-1}-\Sigma_{k+1}^{-1})m_{k+1}+\Sigma_{k+1}^{-1}m_{k+1}=\big{(}\Lambda^{-1}-(1-\tau)\Sigma_{k}^{-1}-\tau\Sigma^{-1}\big{)}m_{k}$ (71)
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+(1-\tau)\Sigma_{k}^{-1}m_{k}+\tau\Sigma^{-1}m$
@@
$\iff m_{k+1}=(I_{d}-\tau\Lambda\Sigma^{-1})m_{k}+\tau\Lambda\Sigma^{-1}m.$
We note that it is the same scheme as for the forward Euler method in the forward-backward scheme. The entropy does not affect the convergence towards the mean, which can be done simply by (preconditioned) gradient descent.
For the variance part, we get
@@
Now, we suppose that all matrices commute,color=violet!30,size=] PCA: artifice de calcul ? Juste pour se simplifier la vie ?Ce calcul fait penser à Woodbury https://en.wikipedia.org/wiki/Woodbury_matrix_identity then
$\Lambda^{-2}\Sigma_{k+1}-2\Lambda^{-1}+\Sigma_{k+1}^{-1}=(\Lambda^{-1}-\tau\Sigma^{-1})^{2}\Sigma_{k}-2(1-\tau)\Lambda^{-1}+2\tau(1-\tau)\Sigma^{-1}$ (73)
$\qquad\qquad\qquad\qquad\qquad\qquad+(1-\tau)^{2}\Sigma_{k}^{-1}$
@@
then the equation is equivalent with

### v2

@@ D.1 Optimal transport maps for mirror descent
Let $\phi:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$ be a strictly convex functional along all acceleration-free curves $\mu_{t}=(\mathrm{T}_{t})_{\#}\mu$, $t\in[0,1]$ with $\mathrm{T}_{t}=(1-t)\mathrm{S}+t\mathrm{T}$ for $\mathrm{T},\mathrm{S}\in L^{2}(\mu)$. Denote for $\mu\in L^{2}(\mu)$, $\phi_{\mu}(\mathrm{T})=\phi(\mathrm{T}_{\#}\mu)$. Since $\phi$ is strictly convex along all acceleration-free curves, by Proposition 13, for all $\mathrm{T}\neq\mathrm{S}\in L^{2}(\mu)$, $\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S})>0$ and thus $\phi_{\mu}$ is strictly convex. Indeed, recall that
@@
1. Observe that problem (42) is equivalent to
@@
We continue this section with additional results relative to the invertibility of mirror maps, which are required in Proposition 2. For a potential energy $\mathcal{V}(\mu)=\int V\mathrm{d}\mu$, since $\nabla_{\mathrm{W}_{2}}\mathcal{V}=\nabla V$, then $\nabla_{\mathrm{W}_{2}}\mathcal{V}$ is invertible provided $\nabla V$ is. This is the case e.g. for $V$ strictly convex. We now state in the two next lemmas conditions for an interaction energy and for the negative entropy to satisfy the invertibility requirements.
@@
Let $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and let $W:\mathbb{R}^{d}\to\mathbb{R}$ be even, $\epsilon$-strongly convex for $\epsilon>0$ and differentiable. Then, for $\mathcal{W}(\mu)=\iint W(x-y)\ \mathrm{d}\mu(x)\mathrm{d}\mu(y)$, $\nabla_{\mathrm{W}_{2}}\mathcal{W}(\mu)$ is invertible.
@@
On one hand, $\nabla_{\mathrm{W}_{2}}\mathcal{W}(\mu)=\nabla W\star\mu$. Moreover, $W$ $\epsilon$-strongly convex is equivalent to
@@
Let $\mu\in\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$ such that its density is of the form $\rho\propto e^{-V}$ with $V:\mathbb{R}^{d}\to\mathbb{R}$ $\epsilon$-strongly convex for $\epsilon>0$. Then, for $\mathcal{H}(\mu)=\int\log\big(\rho(x)\big{missing})\ \mathrm{d}\mu(x)$ with $\rho$ the density of $\mu$ w.r.t the Lebesgue measure, $\nabla_{\mathrm{W}_{2}}\mathcal{H}(\mu)$ is invertible.
@@
The OT problem (42) recovers other OT costs for specific choices of $\phi$. For instance, for $\phi_{\mu}(\mathrm{T})=\frac{1}{2}\|\mathrm{T}\|_{L^{2}(\mu)}^{2}$, it coincides with the squared Wasserstein-2 distance. And more generally, for $\phi_{\mu}^{V}(\mathrm{T})=\int V\circ\mathrm{T}\ \mathrm{d}\mu$, since by Lemma 31, for all $\mathrm{T},\mathrm{S}\in L^{2}(\mu)$,
@@ D.2 Continuous formulation
Let $\phi:L^{2}(\mu)\to\mathbb{R}$ be pushforward compatible and superlinear. Introducing the (mirror) map $\varphi(\mu)=\nabla_{\mathrm{W}_{2}}\phi(\mu)$, we can write informally the mirror descent scheme (4) and its continuous-time counterpart when $\tau\to 0$ as
@@
However, $\frac{\mathrm{d}}{\mathrm{d}t}\varphi(\mu_{t})=\frac{\mathrm{d}}{\mathrm{d}t}\nabla_{\mathrm{W}_{2}}\phi(\mu_{t})=\mathrm{H}\phi_{\mu_{t}}(v_{t})$ where $\mathrm{H}\phi_{\mu_{t}}:L^{2}(\mu_{t})\to L^{2}(\mu_{t})$ is the Hessian operator (defined in Section C.2) such that $\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\phi(\mu_{t})=\langle\mathrm{H}\phi_{\mu_{t}}(v_{t}),v_{t}\rangle_{L^{2}(\mu_{t})}$ and $v_{t}\in L^{2}(\mu_{t})$ is a velocity field satisfying $\partial_{t}\mu_{t}+\mathrm{div}(\mu_{t}v_{t})=0$. Thus, the continuity equation corresponding to the Mirror Flow is given by
@@ D.3 Derivation in specific settings
In this section, we analyze several novel mirror schemes obtained through the use of different Bregman potential maps in (4), and used in various applications in Section 5. We start by discussing the scheme with an interaction energy as Bregman potential. Next, we study mirror descent with negative entropy or KL divergence as Bregman potential. For the last two, we derive closed-forms for the case where every distribution is Gaussian, which is equivalent to working on the Bures-Wasserstein space, and to use the gradient on the Bures-Wasserstein space [43]. In particular, this space is a submanifold of $\mathcal{P}_{2,\mathrm{ac}}(\mathbb{R}^{d})$ and the tangent space is the space of affine maps with symmetric linear term, i.e. of the form $T(x)=b+S(x-m)$ with $S\in S_{d}(\mathbb{R})$.
@@
Consider as Bregman potential an interaction energy $\phi_{\mu}(\mathrm{T})=\frac{1}{2}\iint W\big{(}(T(x)-T(x^{\prime})\big{)}\ \mathrm{d}\mu(x)\mathrm{d}\mu(x^{\prime})$. The mirror descent scheme (4) is given by
@@
For the particular case $W(x)=\frac{1}{2}\|x\|_{2}^{2}$, the scheme can be made more explicit as $\nabla W\star\mu(x)=\int\nabla W(x-y)\ \mathrm{d}\mu(y)=\int(x-y)\ \mathrm{d}\mu(y)=x-m(\mu)$ with $m(\mu)=\int y\mathrm{d}\mu(y)$ the expectation, and thus (53) translates as
$\forall k\geq 0,\ x_{k+1}-m(\mu_{k+1})=x_{k}-m(\mu_{k})-\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}),\;x_{k}\sim\mu_{k}.$ (54)
@@
since $\nabla^{2}W=I_{d}$. On the other hand, the mirror descent scheme (54) can be written as, for all $k\geq 0$,
$y_{k+1}=y_{k}-\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k})(x_{k}),\;y_{k}=x_{k}-m(\mu_{k}),\;x_{k}\sim\mu_{k}.$ (56)
@@
$\frac{\mathrm{d}y_{t}}{\mathrm{d}t}=-\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{t})(x_{t}),\;y_{t}=x_{t}-m(\mu_{t}),\;x_{t}\sim\mu_{t},$ (57)
where $\frac{\mathrm{d}y_{t}}{\mathrm{d}t}=\frac{\mathrm{d}x_{t}}{\mathrm{d}t}-\frac{\mathrm{d}m(\mu_{t})}{\mathrm{d}t}$. Now, by setting $v_{t}(x)=\frac{\mathrm{d}x_{t}}{\mathrm{d}t}$, by integration by part, we have
$\frac{\mathrm{d}}{\mathrm{d}t}m(\mu_{t})=\int x\ \partial_{t}\mu_{t}=-\int x\cdot\mathrm{div}(\mu_{t}v_{t})=\int v_{t}(y)\ \mathrm{d}\mu_{t}(y).$ (58)
Combining the latter equation with (55), we obtain as expected that $\frac{\mathrm{d}y_{t}}{\mathrm{d}t}=\mathrm{H}\phi_{\mu_{t}}[v_{t}](x)$.
@@
Consider the negative entropy $\phi(\mu)=\int\log\big(\rho(x)\big{missing})\ \mathrm{d}\mu(x)$ where $\mathrm{d}\mu(x)=\rho(x)\mathrm{d}x$ and $\phi_{\mu}(\mathrm{T})=\phi(\mathrm{T}_{\#}\mu)$. For such Bregman potential, the mirror scheme (4) can be written for all $k\geq 0$ as
@@
In general, this scheme is not tractable. Nonetheless, supposing that $\mu_{k}=\mathcal{N}(m_{k},\Sigma_{k})$ for all $k\geq 0$, the scheme translates as
$-\Sigma_{k+1}^{-1}(\mathrm{T}_{k+1}(x_{k})-m_{k+1})=-\Sigma_{k}^{-1}(x_{k}-m_{k})-\tau\nabla_{\mathrm{W}_{2}}\mathcal{F}(\mu_{k}),\;x_{k}\sim\mu_{k}.$ (60)
$\bullet$ For an objective functional $\mathcal{F}(\mu)=\mathcal{H}(\mu)+\mathcal{V}(\mu)$ with $V(x)=\frac{1}{2}x^{T}\Sigma^{-1}x$, the scheme is
@@
Assuming $m_{k}=0$ for all $k$, we obtain the following update rule for the covariance matrices:
@@
$\bullet$ For $\mathcal{F}(\mu)=\mathcal{H}(\mu)$, we obtain
$-\Sigma_{k+1}^{-1}(\mathrm{T}_{k+1}(x_{k})-m_{k+1})=-(1-\tau)\Sigma_{k}^{-1}(x_{k}-m_{k}),\;x_{k}\sim\mu_{k}.$ (63)
Assuming $m_{k}=0$ for all $k$, for $\tau<1$, we obtain the following update rule for the covariance matrices:
[…]
$\Sigma_{k+1}$ $=\frac{1}{(1-\tau)^{2}}\Sigma_{k}=\frac{1}{(1-\tau)^{2k}}\Sigma_{0}\underset{\tau\to 0}{\sim}e^{2\tau k}\Sigma_{0}.$ (65)
The continuous time analog of this scheme is thus $\mu_{t}:t\mapsto\mathcal{N}(0,\Sigma_{t})$ with $\Sigma_{t}=e^{2t}\Sigma_{0}$ and the negative entropy decreases along this curve as
$\mathcal{H}(\mu_{t})$ $=\int\log\big(\rho_{t}(x)\big{missing})\ \mathrm{d}\mu_{t}(x)$ (66)
$=\int\log\left(\frac{1}{(2\pi)^{\frac{d}{2}}\sqrt{\det\Sigma_{t}}}e^{-\frac{1}{2}x^{T}\Sigma_{t}^{-1}x}\right)\ \mathrm{d}\mu_{t}(x)$
$=-\frac{d}{2}\log(2\pi)-\frac{1}{2}\log\det(e^{2t}\Sigma_{0})-\frac{1}{2}\mathrm{Tr}\left(\Sigma_{t}^{-1}\int xx^{T}\mathrm{d}\mu_{t}(x)\right)$
$=-\frac{d}{2}\log(2\pi e)-dt-\frac{1}{2}\sum_{i=1}^{d}\log(\lambda_{i}),$
where $(\lambda_{i})_{i}$ denote the eigenvalues of $\Sigma_{0}$. This is much faster than the heat flow for which the negative entropy decreases as [129, Appendix E.2]
@@
with the scheme given by [129, Example 6]
@@
With our notations, the heat flow is the continuous time limit of the scheme (4) for the same objective $\mathcal{F}$ but for a quadratic Bregman potential $\phi_{\mu}(\mathrm{T})=\frac{1}{2}\|\mathrm{T}\|_{L^{2}(\mu)}^{2}$ (which recovers the Wasserstein-2 geometry, hence Wasserstein-2 gradient flows).
@@
Suppose we want to optimize the KL divergence, i.e. a functional of the form $\mathcal{F}(\mu)=\mathcal{G}(\mu)+\mathcal{H}(\mu)$ where $\mathcal{G}(\mu)=\int U\mathrm{d}\mu$. Then, a natural choice of Bregman potential is also a functional of the form $\phi(\mu)=\Psi(\mu)+\mathcal{H}(\mu)$ with $\Psi(\mu)=\int V\mathrm{d}\mu$, with $U$ $\alpha$-convex and $\beta$-smooth relative to $V$.
In that case, we obtain the smoothness of $\mathcal{F}$ relative to $\phi$. Recall we denote ~$\tilde{\mathcal{F}}_{\mu}(\mathrm{T})=\mathcal{F}(\mathrm{T}_{\#}\mu)$ for $\mathrm{T}\in L^{2}(\mu)$. Then for all $\mathrm{T},\mathrm{S}\in L^{2}(\mu)$, we have $\alpha\mathrm{d}_{\tilde{\Psi}_{\mu}}(\mathrm{T},\mathrm{S})\leq\mathrm{d}_{\tilde{\mathcal{G}}_{\mu}}(\mathrm{T},\mathrm{S})\leq\beta\mathrm{d}_{\tilde{\Psi}_{\mu}}(\mathrm{T},\mathrm{S})$, hence
$\mathrm{d}_{\tilde{\mathcal{F}}_{\mu}}(\mathrm{T},\mathrm{S})=\mathrm{d}_{\tilde{\mathcal{H}}_{\mu}}(\mathrm{T},\mathrm{S})+\mathrm{d}_{\tilde{\mathcal{G}}_{\mu}}(\mathrm{T},\mathrm{S})\leq\mathrm{d}_{\tilde{\mathcal{H}}_{\mu}}(\mathrm{T},\mathrm{S})+\beta\mathrm{d}_{\tilde{\Psi}_{\mu}}(\mathrm{T},\mathrm{S})\leq\mathrm{max}(1,\beta)\mathrm{d}_{\phi_{\mu}}(\mathrm{T},\mathrm{S}).$ (69)
@@
We now focus on the case where all measures are Gaussian in order to be able to compute a closed-form, i.e. $U(x)=\frac{1}{2}(x-m)^{T}\Sigma^{-1}(x-m)$, $V(x)=\frac{1}{2}x^{T}\Lambda^{-1}x$ and for all $k\geq 0$, $\mu_{k}=\mathcal{N}(m_{k},\Sigma_{k})$. In this case, recall that $\nabla\log\mu_{k}(x)=-\Sigma_{k}^{-1}(x-m_{k})$. Then, at each step, the mirror descent scheme (4) writes for $x_{k}\sim\mu_{k},\;k\geq 0$ as
$\nabla V(x_{k+1})+\nabla\log\big(\mu_{k+1}(x_{k+1})\big{missing})=\nabla V(x_{k})+\nabla\log\big(\mu_{k}(x_{k})\big{missing})-\tau\big{(}\nabla U(x_{k})+\nabla\log\big(\mu_{k}(x_{k})\big{missing})$
@@
$\quad\quad\quad=\big{(}\Lambda^{-1}-(1-\tau)\Sigma_{k}^{-1}-\tau\Sigma^{-1}\big{)}x_{k}+(1-\tau)\Sigma_{k}^{-1}m_{k}+\tau\Sigma^{-1}m.$ (70)
@@
$(\Lambda^{-1}-\Sigma_{k+1}^{-1})m_{k+1}+\Sigma_{k+1}^{-1}m_{k+1}=\big{(}\Lambda^{-1}-(1-\tau)\Sigma_{k}^{-1}-\tau\Sigma^{-1}\big{)}m_{k}(1-\tau)\Sigma_{k}^{-1}m_{k}+\tau\Sigma^{-1}m$
@@
$\iff m_{k+1}=(I_{d}-\tau\Lambda\Sigma^{-1})m_{k}+\tau\Lambda\Sigma^{-1}m.$ (71)
We note that the latter update on the means coincides with the forward Euler method in the forward-backward scheme, see (116) in Appendix F, which uses as Bregman potential $\phi=\Psi$. Thus, the entropy does not affect the convergence towards the mean, which can be done simply by (preconditioned) gradient descent.
For the covariance part, we get
@@
Now, supposing that all matrices commute, we get
$\Lambda^{-2}\Sigma_{k+1}-2\Lambda^{-1}+\Sigma_{k+1}^{-1}=(\Lambda^{-1}-\tau\Sigma^{-1})^{2}\Sigma_{k}-2(1-\tau)\Lambda^{-1}+2\tau(1-\tau)\Sigma^{-1}$
$\qquad\qquad\qquad\qquad\qquad\qquad+(1-\tau)^{2}\Sigma_{k}^{-1}$ (73)
@@
the update on covariances is equivalent to